What numbers are missing from the pattern below? Enter your answer, using a comma to separate each number.
3, 4, 6, 9, ? , ? , ? , ? , 39, 48, 58, 69

Answer by Fombitz(32388)

The difference increases by 1 at each successive number,
3+1=4
4+2=6
6+3=9
9+4=13
13+5=18
18+6=24
24+7=31
31+8=39
48+9=58
58+10=69